# Whom We Trust, What We Fear: COVID-19 Fear and the Politics of Information

### v1

@@ 7. Conclusion
Building on a foundation of behavioral and demographic data, this study explores the factors driving fear of COVID-19 infection, with a specific focus on how information sources and demographic characteristics influence this fear. We examine how various channels—such as social media, traditional journalism, government briefings, and interpersonal communication—are associated with differences in perceived fear across the U.S. population. By analyzing data spanning May 2021 to June 2022, encompassing the Delta and Omicron waves, we capture a critical period of the pandemic marked by significant informational and emotional turbulence.
@@
Recognizing the potential biases inherent in digital platforms—particularly their skew toward users with consistent internet access—we contextualize our findings within the socio-economic and geographic diversity of the United States. Regional disparities in media habits, internet penetration, and education levels emphasize the importance of a geographically nuanced approach to understanding informational dynamics and public emotion.
Our analysis reveals that the type of information source individuals rely on plays a central role in shaping emotional responses during health crises. The credibility, framing, and volume of content can either amplify or mitigate public fear, with misinformation posing a significant threat to risk perception and trust. This underscores the critical importance of ensuring access to verified, reliable information sources to foster informed decision-making and emotional resilience.
Using causal inference techniques, we assessed the relative influence of age, education, and information sources on perceived fear. Among these factors, the type of information source emerged as the most significant driver. Furthermore, our findings demonstrate that demographic characteristics not only influence emotional reactions but also shape preferences for specific types of information, revealing distinct behavioral patterns across age and education groups. Taken together, these insights provide critical guidance for designing more targeted and trustworthy communication strategies in future public health emergencies, where managing both the transmission of disease and the spread of fear is essential.
@@
Appendix A Extended analysis on clustering
@@
In addition to the analysis presented in Section 6—specifically in the U.S. states subsection—, we provide additional details on clustering using $\hat{\sigma}_{s^{\prime}}$. We also repeat the analysis with alternative inputs, including $\phi_{s^{\prime}}$ alone and in combination with $\hat{\sigma}_{s^{\prime}}$.
Figure 10. Clusters derived from K-means analysis based on average reliance on each singleton information source, $\hat{\sigma}_{s^{\prime}}$, across U.S. states, stratifying by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections.
@@
Clustering using $\hat{\sigma}_{s^{\prime}}$. Alongside the results shown in Section 6, we incorporate demographic information by computing, for each U.S. state, the average reliance on each singleton information source, grouped by Age, resulting in 27 values per state. As shown in Figure 10, this adjustment reduced misclassifications in the 2024 elections, with only Wisconsin and Alaska assigned to the opposite cluster—but led to an increase in misclassifications for the 2020 elections. A similar improvement is observed when combining information source preferences with Education levels.
@@
Clustering using $\phi_{s^{\prime}}$. For each U.S. state, we compute the average fear score for each singleton information source, $\phi_{s^{\prime}}$, without stratifying by Age or Education, resulting in nine values per state. Figure 11 presents the resulting clusters and compares them with the outcomes of the 2020 and 2024 political elections.
Compared to the clusters obtained using $\hat{\sigma}_{s^{\prime}}$, a larger number of states are misclassified in both elections—specifically, Colorado, Connecticut, Delaware, Illinois, Maine, Massachusetts, Minnesota, New Hampshire, Oregon, Rhode Island, Texas, Vermont, Virginia, and Washington. Notably, swing states are misclassified only in the 2020 elections, but appear in the correct cluster for 2024. Additionally, Alaska and Hawaii are now correctly classified in both elections.
@@
Figure 11. Clusters derived from K-means analysis based on the average respondents’ fear score for each singleton information source, $\phi_{s^{\prime}}$, across U.S. states, without stratification by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections.
@@
Clustering using $\hat{\sigma}_{s^{\prime}}$ and $\phi_{s^{\prime}}$. For each U.S. state, we compute the average reliance $\hat{\sigma}_{s^{\prime}}$ and fear scores $\phi_{s^{\prime}}$ for each singleton information source without stratifying by Age or Education, resulting in 18 values per state. Figure 12 presents the resulting clusters obtained and compares them to the outcomes of the 2020 and 2024 U.S. political elections.
@@
Figure 12. Clusters derived from K-means analysis based on the average reliance $\hat{\sigma}_{s^{\prime}}$ on each singleton information source and average respondents’ fear score $\phi_{s^{\prime}}$ for them across U.S. states, without stratification by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections.
Appendix B Ethical declaration
The Ethics Board (IRB) of IMDEA Networks Institute gave ethical approval for this work on 2021/07/05. IMDEA Networks has signed Data Use Agreements with Facebook, Carnegie Mellon University (CMU) and the University of Maryland (UMD) to access their data, specifically UMD project 1587016-3 entitled C-SPEC: Symptom Survey: COVID-19 and CMU project STUDY2020_00000162 entitled ILI Community-Surveillance Study. The data used in this study was collected by the University of Maryland through The University of Maryland Social Data Science Center Global COVID-19 Trends and Impact Survey in partnership with Facebook. Informed consent has been obtained from all participants in this survey by this institution. All the methods in this study have been carried out in accordance with relevant ethical and privacy guidelines and regulations.
Appendix C Availability of materials and data
All datasets used in this work are open-source and publicly available. In particular, we used official surveillance data from the COVID-19 Data Hub (Guidotti and Ardia, 2020; Guidotti, 2022) and additional data from the Delphi US CTIS (Delphi Group, CMU, 2020) on demographic characteristics (such as age and education), sources of COVID-19-related information, and reported levels of fear. The microdata of the Delphi US CTIS survey from which the aggregated data was obtained cannot be shared, as per the Data Use Agreements signed with Facebook, Carnegie Mellon University (CMU), and the University of Maryland (UMD). Thus, the authors do not have permission to share the microdata of the Delphi US CTIS survey. The Delphi US CTIS survey data (in aggregated form) and the code are openly accessible at https://github.com/paolocastagno/covid19-fear.
Appendix D Declaration of competing interest
@@
Appendix E Funding
This work was supported by grants from TED2021-131264B-I00 (SocialProbing) funded by MCIN/AEI/10.13039/501100011033 and the European Union “NextGenerationEU”/PRTR.

### v2

@@ Conclusion
Building on behavioral and demographic data, this study explores the factors driving fear of COVID-19 infection, focusing on how information sources and demographics shape emotional responses. We examine how channels such as social media, traditional journalism, government briefings, and interpersonal communication relate to differences in perceived fear across the U.S., analyzing data from May 2021 to June 2022, covering the Delta and Omicron waves.
@@
Recognizing potential biases in digital platforms, particularly toward users with consistent internet access, we contextualize our findings within the socio-economic and geographic diversity of the U.S. Regional disparities in media habits, internet penetration, and education underscore the importance of a geographically nuanced approach to understanding information dynamics and public emotion. Moreover, this study analyzes data exclusively from the U.S. Delphi CTIS survey, which is restricted to American respondents. While this enables detailed causal inference within the U.S. context, the work could be extended to other countries if comparable longitudinal panel data on COVID-19 fear, information sources, and related behaviors become available in the future.
Our analysis shows that COVID-19–related fear associated with the considered information sources is strongly correlated across sources and with the number of confirmed infections. Using causal inference techniques, we quantified the effects of age, education, and information sources on perceived fear. Information sources emerged as the most significant driver, while demographics also shape emotional reactions and preferences for specific types of information, revealing distinct behavioral patterns across age and education groups. These results underscore the pivotal role of the information environment in shaping public fear during health crises. The credibility, framing, and volume of content can amplify or mitigate fear, with misinformation posing a serious threat to risk perception and trust. Ensuring access to verified, reliable sources is therefore critical for informed decision-making and emotional resilience.
@@
Checklist
  1. 1.
    For most authors…
    1. (a)
      Would answering this research question advance science without violating social contracts, such as violating privacy norms, perpetuating unfair profiling, exacerbating the socio-economic divide, or implying disrespect to societies or cultures? Yes, see the Data and Ethical declaration sections.
    2. (b)
      Do your main claims in the abstract and introduction accurately reflect the paper’s contributions and scope? Yes, the Abstract and the Introduction section state the main contributions of the paper, and we explicitly summarize them in a dedicated Contributions subsection within the Introduction section.
    3. (c)
      Do you clarify how the proposed methodological approach is appropriate for the claims made? Yes, in the Methods section.
    4. (d)
      Do you clarify what are possible artifacts in the data used, given population-specific distributions? Yes, in the Data and Conclusion sections.
    5. (e)
      Did you describe the limitations of your work? Yes, in the Data, Methods, an Results sections.
    6. (f)
      Did you discuss any potential negative societal impacts of your work? NA.
    7. (g)
      Did you discuss any potential misuse of your work? NA.
    8. (h)
      Did you describe steps taken to prevent or mitigate potential negative outcomes of the research, such as data and model documentation, data anonymization, responsible release, access control, and the reproducibility of findings? NA.
    9. (i)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? Yes.
  2. 2.
    Additionally, if your study involves hypotheses testing…
    1. (a)
      Did you clearly state the assumptions underlying all theoretical results? NA
    2. (b)
      Have you provided justifications for all theoretical results? NA
    3. (c)
      Did you discuss competing hypotheses or theories that might challenge or complement your theoretical results? NA
    4. (d)
      Have you considered alternative mechanisms or explanations that might account for the same outcomes observed in your study? NA
    5. (e)
      Did you address potential biases or limitations in your theoretical framework? NA
    6. (f)
      Have you related your theoretical results to the existing literature in social science? NA
    7. (g)
      Did you discuss the implications of your theoretical results for policy, practice, or further research in the social science domain? Yes, in the Conclusion section.
  3. 3.
    Additionally, if you are including theoretical proofs…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? NA
    2. (b)
      Did you include complete proofs of all theoretical results? NA
  4. 4.
    Additionally, if you ran machine learning experiments…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? Yes, see the Availability of materials and data section.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? NA
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? NA
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? NA
    5. (e)
      Do you justify how the proposed evaluation is sufficient and appropriate to the claims made? NA
    6. (f)
      Do you discuss what is “the cost“ of misclassification and fault (in)tolerance? NA
  5. 5.
    Additionally, if you are using existing assets (e.g., code, data, models) or curating/releasing new assets, without compromising anonymity
    1. (a)
      If your work uses existing assets, did you cite the creators? Yes, we have cited relevant papers.
    2. (b)
      Did you mention the license of the assets? NA
    3. (c)
      Did you include any new assets in the supplemental material or as a URL? NA
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? Yes, see the Ethical declaration section.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? Yes, the data we used does not contain personally identifiable information or offensive content.
    6. (f)
      If you are curating or releasing new datasets, did you discuss how you intend to make your datasets FAIR? NA
    7. (g)
      If you are curating or releasing new datasets, did you create a Datasheet for the Dataset? NA
  6. 6.
    Additionally, if you used crowdsourcing or conducted research with human subjects, without compromising anonymity
    1. (a)
      Did you include the full text of instructions given to participants and screenshots? NA
    2. (b)
      Did you describe any potential participant risks, with mentions of Institutional Review Board (IRB) approvals? NA
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? NA
    4. (d)
      Did you discuss how data is stored, shared, and deidentified? NA
Appendix A: Extended analysis on clustering
@@
In addition to the analysis presented in the Results section—specifically in the Geographical distribution of sources: U.S. states subsection—, we provide additional details on clustering using $\hat{\sigma}_{s^{\prime}}$. We also repeat the analysis with alternative inputs, including $\phi_{s^{\prime}}$ alone and in combination with $\hat{\sigma}_{s^{\prime}}$. In each case, we present the clusters obtained using K-means, noting that similar results were obtained with hierarchical and spectral clustering methods.
@@
Clustering using $\hat{\sigma}_{s^{\prime}}$. Alongside the results shown in the Results section, we incorporate demographic information by computing, for each U.S. state, the average reliance on each singleton information source, grouped by Age, resulting in 27 values per state. As shown in Figure 11, this adjustment reduced misclassifications in the 2024 elections, with only Wisconsin and Alaska assigned to the opposite cluster—but it led to a slight increase in misclassifications for the 2020 elections. A similar improvement is observed when combining information source preferences with Education levels.
Figure 11: Clusters derived from K-means analysis based on average reliance on each singleton information source, $\hat{\sigma}_{s^{\prime}}$, across U.S. states, stratifying by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections.
@@
Clustering using $\phi_{s^{\prime}}$. For each U.S. state, we compute the fear score for each singleton information source, $\phi_{s^{\prime}}$, without stratifying by Age or Education, resulting in nine values per state. Figure 12 presents the resulting clusters and compares them with the outcomes of the 2020 and 2024 political elections.
Compared to the clusters obtained using $\hat{\sigma}_{s^{\prime}}$, a larger number of states are misclassified in both elections—specifically, Colorado, Connecticut, Delaware, Illinois, Maine, Massachusetts, Minnesota, New Hampshire, Oregon, Rhode Island, Texas, Vermont, Virginia, and Washington. Notably, swing states are misclassified only in the 2020 elections, but appear in the correct cluster for 2024. Additionally, Alaska is now correctly classified in both elections.
@@
Figure 12: Clusters derived from K-means analysis based on the average respondents’ fear score for each singleton information source, $\phi_{s^{\prime}}$, across U.S. states, without stratification by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections.
@@
Clustering using $\hat{\sigma}_{s^{\prime}}$ and $\phi_{s^{\prime}}$.
Figure 13: Clusters derived from K-means analysis based on the average reliance $\hat{\sigma}_{s^{\prime}}$ on each singleton information source and average respondents’ fear score $\phi_{s^{\prime}}$ for them across U.S. states, without stratification by Age or Education. These clusters are compared with the results of the a) 2020 and b) 2024 U.S. political elections.
For each U.S. state, we compute the average reliance $\hat{\sigma}_{s^{\prime}}$ and fear scores $\phi_{s^{\prime}}$ for each singleton information source without stratifying by Age or Education, resulting in 18 values per state. Figure 13 presents the resulting clusters obtained and compares them to the outcomes of the 2020 and 2024 U.S. political elections.
@@
Appendix A Ethical declaration
Further details will be provided upon acceptance, in accordance with double-blinded submission requirements.
Appendix B Availability of materials and data
All datasets used in this work will be publicly available. Further details will be provided upon acceptance, in accordance with double-blinded submission requirements.
Appendix C Declaration of competing interest
@@
Appendix D Funding
Further details will be provided upon acceptance, in accordance with double-blinded submission requirements.